40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 4 20:54:40 2020, gautam, Update, SUS, MC1 Slow Bias issues 
    Reply  Thu Jun 4 21:32:00 2020, Koji, Update, SUS, MC1 Slow Bias issues 
       Reply  Thu Jun 25 15:11:00 2020, gautam, Update, SUS, MC1 coil driver resistance quartered MCF.pdf
          Reply  Sun Jun 28 15:30:52 2020, gautam, Update, SUS, MC1 sat-box de-lidded 20200628T144138.jpg
             Reply  Sun Jun 28 16:29:58 2020, rana, Update, SUS, MC1 sat-box de-lidded 
                Reply  Sun Jun 28 17:36:35 2020, gautam, Update, SUS, MC1 sat-box de-lidded 
                   Reply  Mon Jun 29 11:55:46 2020, gautam, Update, SUS, MC1 sat-box de-lidded SatBoxSwap.jpg
                      Reply  Mon Jun 29 20:30:53 2020, Koji, Update, SUS, MC1 sat-box de-lidded 
Message ID: 15431     Entry time: Thu Jun 25 15:11:00 2020     In reply to: 15377     Reply to this: 15434
Author: gautam 
Type: Update 
Category: SUS 
Subject: MC1 coil driver resistance quartered 

I implemented this change today. We only had 100 ohm, 3W resistors in stock (no 200 ohm with adequate power rating). Assuming 10 V is dropped across this resistor, the power dissipation is V^2/R ~ 1 W, so we should have sufficient margin. DCC entry has been updated with new schematic and photo of the component side of the board. Note that the series resistance of the fast actuation path was untouched.

As expected, the requested voltage no longer exceeds the Acromag DAC range, it is now more like 2.5 V. However, I still notice that the MC REFL spot moves somewhat diagonally on the camera image - so maybe the coil gains are seriously imbalanced? Anyway, the WFS control signals can once again be safely offloaded to the slow bias voltages once again, preserving the fast ADC range for other actuation.

The Johnson noise of the series resistor has now increased by a factor of 2, from ~6.4 pA/rtHz to 12.8 pA/rtHz. Assuming a current to force coefficient of 1.6 mN/A per coil, the length noise of the cavity is expected to be 12.8e-12 * 0.064/0.25/(2*pi*100)^2 ~ 8e-18 m/rtHz at 100 Hz. In frequency units, this is 80 uHz/rtHz. I think our IMC noise is at least 10 times higher than this at 100 Hz (in any case, the noise of the coil driver is NOT dominated by the series resistance). Attachment #1 confirms that there isn't any significant MCF noise increase, and I will check with the arm cavity too. Nevertheless, we should, if possible, align the optic better and use as high a series resistance as possible.

The watchdog for MC1 was disabled and the board was pulled out for this work. After it was replaced, the IMC re-locks readily.

Quote:

But this does not solve the MC1 issue. Only we can do right now is to make the output resister half, for example.

Attachment 1: MCF.pdf  41 kB  Uploaded Thu Jun 25 16:34:02 2020  | Hide | Hide all
MCF.pdf
ELOG V3.1.3-